OpenStudy (anonymous):

1/6x=x+6/2x^2+1/3x^2

OpenStudy (anonymous):

What is it asking you to do?

OpenStudy (anonymous):

do you need to find X?

OpenStudy (anonymous):

Solve each equation remember to check for extraneous solutions

OpenStudy (anonymous):

If you're solving for x then the answer would be: \[x = 0, -\frac{ 1 }{ 4 }\]

OpenStudy (anonymous):

1/2x-1/4x^2=2/x^2

OpenStudy (anonymous):

Could you show more work please?